# A Visual Case Study of the Training Dynamics in Neural Networks

### v1

@@ 4.2 Transfer Experiments
Among the lessons learned from training very large models is the importance of careful data engineering (Touvron et al., 2023b; Team, 2024). Cabannes et al. (2024a) has highlighted that certain sources of data may facilitate the learning of invariances, while Abbe et al. (2023) discusses how data curation enables models to escape saddle points more quickly. These insights are consistent with the observations we made regarding our problem. In the sparse modular addition problem, the number of unique sequences is equal to $p^{T}$, which increases rapidly with both $p$ and $T$ and makes the problem quite hard to learn. In particular, when setting $T=12$ and limiting the training set to $n=2048$ data points, training for 1000 epochs does not result in any learning for $p>3$.
@@
As previously seen in Figure 4, when training with $p=2$, we often find circuits that capture both permutation and suffix invariants. These invariants generalize for any $p\in{\mathbb{N}}$ when $k$ and $T$ are fixed. Consequently, initializing models with these invariants makes learning the sparse modular addition problem much easier. This observation was made after conducting the following experiments: we first trained a model with sequences in ${\mathbb{F}}_{p}$ for $p=2$ over 1000 epochs, before switching the dataset to sequences in ${\mathbb{F}}_{p}$ for $p=3$ for another 1000 epochs. We found that this procedure significantly facilitates the learning of the sparse modular addition problem for $p=3$, which we summarize in Figure 9 where each bar plot is obtained by over $250$ runs. Remarkably, we found that the only models that achieved 100% test accuracy were those that captured both the token and permutation invariances after the first 1000 epochs. Specifically, these were the models that created six sequence embedding clusters, as shown in Figures 4 and 4, rather than those depicted in Figures 4 or 4.
Figure 11 shows the final circuit found in one of our finetuning experiments. The training was initialized with the circuit in Figure 4, after adding a token embedding to encode for $x=2$, resulting in Figure 11. The final embeddings are not that far from the initial one, with the transformer having learned to mainly pay attention to the non-spurious tokens that are not equal to 2. It also pays some attention to 0 and 1 in positions $t=7$ and $t=10$. However, this faulty attention is corrected by the value matrix. Once again, the final configuration seems somewhat close to the initial one, as shown by the attention pattern reported in Figure 8. This is consistent with the observations made in the previous subsections.
Figure 10: Same circuit as in Figure 4 with an additional token embedding for $p=3$.
Figure 11: Circuit learned after $1000$ epochs of finetuning with $p=3$.
@@
4.3 Loss Spikes
One interesting aspect of our sandbox is that it generates loss spikes that we can study quite precisely. Our visual inspection showcases two aspects beyond loss spikes. The high curvature of multi-layer perceptrons with heavy weights, or with numerous small correlated weights, as well as the high curvature of the RMS normalization layer near the origin.
Figure 12: Loss spikes are linked to the high curvature of internal network functions. A small update to an element $x$ can result in a substantial change to its normalized version $x/\left\|x\right\|$, significantly altering the network’s subsequent behavior.
Figure 13: Loss spikes (both left) resulting from a small change from one iteration to another in sequence embeddings that are close to the decision boundaries of the subsequent feedforward layer (both right).
Figure 12 illustrates that a small modification to an element can result in a disproportionately large change in its normalized version. In theory, gradient points towards directions that would reduce the training loss. However, considering a large step size in these directions could be counterproductive. This is especially true for functions with high curvature, such as the normalization layer $f(x)=x/\left\|x\right\|$ near the origin. At any point $x_{t}$, the gradient descent update rule suggests that one can update $x_{t+1}$ as $x_{t}-\eta_{t}u_{t}$ without changing $f(x_{t})$, where $u_{t}=x_{t}/\|x_{t}\|$ and $\eta_{t}$ is the learning rate. This holds true only if the learning is small enough, $\eta_{t}<\|x_{t}\|$. When $x_{t}$ is close to zero, ensuring $\eta_{t}<\|x_{t}\|$ becomes challenging, particularly if the step size $\eta_{t}$ was predetermined by some scheduler. This behavior is related to the "edge-of-stability" phenomenon highlighted by Cohen et al. (2022), further theoretical insights being provided by Cabannes et al. (2024b). Interestingly, this analysis suggests removing some loss spikes by smoothing out the normalization layer. For example, consider using $f(x)=\sigma(\left\|x\right\|)x/\left\|x\right\|$ where $\sigma$ is a smooth function with $\sigma(0)=0$ and $\sigma([1,\infty))=\{1\}$. This demonstrates the usefulness of our visual sandbox in gaining insights and building intuition, which can then be validated on a larger scale in separate projects.
To conclude, our visual sandbox enabled us to peek into some reasons behind loss spikes in large neural networks and to build intuition on how to circumvent them, giving us ideas for large-scale experiments, which, if successful, could help increase the amount of intelligence learned for a given amount of training compute. Connection to gradient norms and sparsity is discussed in Appendix A.2 and A.3.
Conclusion
This paper aims to advance our understanding of neural network training dynamics through a detailed study of a small-scale transformer model, facilitated by a novel visual sandbox. This tool allows us to observe and analyze the model’s internal mechanisms vividly, providing both theoretical and practical insights that could lead to more efficient training strategies.
Our findings highlight a hierarchical learning process where low-level features are developed first, followed by a more predictable refinement of the model. We also emphasize the impact of initial configurations on the final model outcomes and the seemingly unpredictable nature of some feature learning phase transitions during training. Additionally, we explored the transferability of learned behaviors, which relates to the importance of data curation and curriculum learning in enhancing model performance. We also addressed loss spikes caused by high curvature in the model’s internal functions and proposed potential solutions to mitigate these issues. Looking ahead, future work will aim to apply these insights to larger and more complex models, assess the scalability of observed phenomena, and enhance our visualization tools for higher-dimensional data.
Societal Impact.
This paper aims to understand the complex dynamics at play in neural networks. The insights gained from our research stream could help in designing more efficient and robust neural networks, ultimately aiding in reducing the computational cost and environmental impact of training large models.
Appendix A Appendix
A.1 Reproducibility Statement
We open-source the code at https://github.com/facebookresearch/pal and provide the implementation details in Section 2.2 to ensure reproducibility.
A.2 Understanding Saddle points in Loss Optimization via Gradient Norms
When the model’s training stagnates, the loss plateau for several epochs before it decreases again. This behavior may be understood by the optimizer reaching a saddle point in the loss function w.r.t. the model’s parameter, that is the gradient of the loss is (almost) zero, but the optimization has not reached yet a local minimum. We illustrate this phenomenon in the full-batch and mini-batch setup, respectively in Figures 16 and 16. The connection between gradient norms and learning phases is salient in Figure 16. We can see that loss drops occur in tandem with high gradient norms for each layer. As studied in Section 4.1, these drops correspond to successive learning phases. This is even more salient in the last subfigure of Figure 16 where the pics precisely match the drops. This shows the connection between the exit of a saddle point and high gradient norms. Similarly, in Figure 16, we can see that the first drop in the loss (around epoch $250$ in the first plot) appears when the first increase in gradient norm occurs and the second point of inflection (around epoch $700$) also match an infection point on the gradient norms. It should be noted that this is less salient compared to the full-batch setup.
@@ A.2 Understanding Saddle points in Loss Optimization via Gradient Norms
Figure 15: Connection between loss spikes and gradient norm in full-batch setup. From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale. We see that the learning phases of Section 4.1 appear in tandem with high gradient norms. This can be seen in the last subfigure where the three pics correspond to the loss drops and their corresponding plateaus
Figure 16: Mini-batch setup. This study, akin to Figure 16, makes the connection between gradient norm and loss profile even more salient for our problem. The first loss drop appears simultaneously with the first increase in gradient norms. Loss spikes occur in tandem with spikes in gradient norms.
However, this setup enables us to study another training behavior mentioned in Sectin 4.3. Indeed, we observe in Figure 16 that the loss spikes appear in tandem with high gradient norms, indicating that when a too-large step size deviates the model from its current small loss region, it is taken back to where it was with large updates. It has been shown in the literature (Ilbert et al., 2024; Foret et al., 2021; Zhang et al., 2024) that studying the gradient and the hessian of the loss could provide valuable insights on the neural network optimization both from a theoretical and experimental point of view. We believe using our visual sandbox could help in future work on such an investigation.
A.3 Impact of Activation Sparsity on Models Performance
Sparsity is a phenomenon of interest in many fields such as signal processing, neuroscience, and machine learning (Mairal et al., 2009; Chen et al., 1998; Barth & Poulet, 2012). Recent studies focused on the sparsity in deep neural network activations. In particular, Li et al. (2023) showed that trained transformers have sparse activations and concluded that it was caused by the training dynamics rather than by a compact representation of the training data as commonly thought in computer vision and NLP. Mirzadeh et al. (2024) observed a similar phenomenon and showed how to leverage sparsity to reduce the inference cost of large language models. Inspired by this line of work, we study the activation sparsity of our model from a performance viewpoint. It should be noted that those works study deep transformers and identified that the sparsity increases with the depth while we only consider a one-layer transformer.
Following the framework from Li et al. (2023), we recall that the activation sparsity corresponds to the percentage of non-zero entries of the feed-forward activation map. Without loss of generality, the feed-forward block is an MLP with weights $W_{1},W_{2}$ and a non-linear activation $\sigma$ that outputs for any input $x$ a vector $z=W_{1}\sigma(W_{2}x)$. Formally, the activation sparsity is the percentage of non-zero entries in $\sigma(W_{2}x)$ and takes values in $[0,1]$. In the classical setting with ReLU activation (Fukushima, 1969), this is equivalent to computing the percentage of non-negative neurons before the activation. However, some activations do not have non-negative outputs. This is the case of the SiLU (Elfwing et al., 2018), used in Llama models (Touvron et al., 2023a), and of the GeLU Hendrycks & Gimpel (2023) used in Falcon (Almazrouei et al., 2023), PaLM (Chowdhery et al., 2024), and in our transformer implementation. Instead of replacing such activations by a ReLU (Mirzadeh et al., 2024; Li et al., 2023), we compute a smoothed sparsity [^1] This is akin to using the $\ell_{1}$-norm, respectively the nuclear norm, instead of the $\ell_{0}$ quasi-norm, respectively the rank (Ilbert et al., 2024; Gribonval & Nielsen, 2003). as the percentage of entries with an absolute value lower than $\varepsilon>0$. A sparsity of $1$ means that all entries are $\varepsilon$-close to 00 (i.e., sparse activations) and a sparsity of 00 means that all entries are at least $\varepsilon$-away from 00 (i.e., dense activations).

### v2

@@ A.3 Transferability of Circuits
Among the lessons learned from training very large models is the importance of careful data engineering (Touvron et al., 2023b; Llama Team, 2024). Cabannes et al. (2024a) has highlighted that certain sources of data may facilitate the learning of invariances, while Abbe et al. (2023) discusses how data curation enables models to escape saddle points more quickly. These insights are consistent with the observations we made regarding our problem. In the sparse modular addition problem, the number of unique sequences is equal to $p^{T}$, which increases rapidly with both $p$ and $T$ and makes the problem quite hard to learn. In particular, when setting $T=12$ and limiting the training set to $n=2048$ data points, training for 1000 epochs does not result in any learning for $p>4$.
@@
Impact on the performance.
As previously seen in Figure 13, when training with $p=2$, we often find circuits that capture both permutation and suffix invariants. These invariants generalize for any $p\in{\mathbb{N}}$ when $k$ and $T$ are fixed. Consequently, initializing models with these invariants makes learning the sparse modular addition problem much easier. This observation was made after conducting the following experiments: we first trained a model with sequences in ${\mathbb{F}}_{p}$ for $p=2$ over 1000 epochs, before switching the dataset to sequences in ${\mathbb{F}}_{p}$ for $p=3$ for another 1000 epochs. We found that this procedure significantly facilitates the learning of the sparse modular addition problem for $p=3$, which we summarize in Figure 14 where each bar plot is obtained by over $250$ runs. Remarkably, we found that the only models that achieved 100% test accuracy were those that captured both the token and permutation invariances after the first 1000 epochs. Specifically, these were the models that created six sequence embedding clusters, as shown in Figures 1 and 13, rather than those depicted in Figures 13 or 13.
Impact on the learned circuits.
Figure 15(b) shows the final circuit found in one of our finetuning experiments. The training was initialized with the circuit in Figure 13, after adding a token embedding to encode for $x=2$, resulting in Figure 15(a). The final embeddings are not that far from the initial one, with the transformer having learned to mainly pay attention to the non-spurious tokens that are not equal to 2. It also pays some attention to 0 and 1 in positions $t=7$ and $t=10$. However, this faulty attention is corrected by the value matrix. Once again, the final configuration seems somewhat close to the initial one, as shown by the attention pattern reported in Figure 8. This is consistent with the observations made in the previous subsections.
(a) Initialization with an additional token embedding for $p=3$.
(b) Circuit learned after $1000$ epochs of finetuning with $p=3$.
Figure 15: Transfer of clustering head. The obtained circuit is akin to Fig. 13 where the value corrects the faulty attention.
Benefits of transfer.
@@
A.4 Impact of the Hyperparameters
In this section, we conduct ablation studies on the training hyperparameters. In Fig. 16(a), we observe the high variability of the test accuracy for a fixed seed after $1000$ epochs when varying the batch size, hidden dimension, learning rates, and MLP learning rates. Fig. 16(b) is similar with more regularity as it is averaged over $100$ runs with different seeds. This study shows that the exit of saddle points is highly dependent on the hyperparameters.
(a) Single run with 1000 epochs.
(b) Averaged over $100$ runs with standard deviations.
Figure 16: Ablation studies regarding test accuracy as a function of batch size, hidden neurons, learning rates, and MLP learning rates discount factor for a single run with 1000 epochs. We ensured consistency in initial weights and batch designs when changing hyperparameters. We observe (mostly in Fig. 16(b)) some regularity in the effect of hyperparameters on the resulting test accuracy.
Appendix B Beyond Low-dimensional Embeddings
Relying on our visual sandbox, we studied several phenomena on the training dynamics of neural networks from the learning of the different parts of the network in Section 4.1 to the loss spikes occurring during the optimization in Section 4.2 through the efficiency of transfer learning in Section A.3. We also analyzed the connection between saddle points, loss spikes, and gradient norms in Sections 4.1 and 4.2 and the impact of the activation sparsity on the models’ performance in Section A.2. The low-dimensional embeddings allowed us to pair loss profiles with the visualization of each layer to decipher the training behaviors. However, we note that the studied behaviors and the insights we obtain can be analyzed independently of the embedding dimension $d$.
@@ Appendix B Beyond Low-dimensional Embeddings
Figure 17: Connection to saddle points. This study is akin to Fig. 3 with $d=3$ in the top plot and $d=64$ in the bottom one. From left to right: Evolution of train and test losses, the corresponding accuracies, the evolution of gradient norms for each layer in log-scale, and the similar evolution in linear scale in full-batch. Akin to Fig. 3, we see that the learning phases studied in Section 4.1 appear in tandem with high gradient norms. This can be seen in the last subfigure where the pics correspond to the loss drops and spikes.
Motivation.
Taking an embedding size $d=2$ was mainly motivated by visualization purposes and we saw that such a setting was already challenging enough to observe training phenomena of interest in more practical settings. For the sake of self-consistency, we extend our experiments with higher-dimensional embeddings and show that our insights remain valid in more practical settings with $d>2$.
B.1 Two-Phase Learning and Saddle Points
In this section, we extend the experiments of Section 4.1 for $d>2$. In Fig. 17, we consider $d=3$ (top) and $d=64$ (bottom). We obtain similar conclusions than in the case $d=2$ in Fig. 3 where the loss drops occur in tandem with high gradient norms for each layer, hinting at the exit of saddle points.
B.2 Loss Spikes and Gradient Norms
In this section, we extend the experiments of Section 4.2 for $d>2$. In Fig. 18, we consider $d=3$ (top) and $d=64$ (bottom). We observe a similar behavior as in the case $d=2$ in Fig. 7 where the loss spikes and gradient norm spikes occur at the same time. This again highlights the connection between gradient norms and loss spikes.
Figure 18: Loss spikes and gradient norms. This study is akin to Fig. 7 in mini-batch with $d=3$ in the top plot and $d=64$ in the bottom one. Again, we observe that the loss spikes and gradient norms spikes appear in tandem.
B.3 Activation Sparsity and Performance
Finally, we extend the experiments of Apppendix A.2 with an embedding size $d\in\{2,3,4,8,16,32\}$ in Fig. 19. We first note that the higher the embedding size, the more the model succeeds at the task. Especially, as of $d=8$, all the models are successful, i.e., they all achieve an accuracy higher than $0.9$ (as defined in Section A.2). This is expected given that imposing low-dimensional embeddings limits the expressiveness and the generalization power of our model. It should be noted that this was one of the many challenges of our study: obtaining a generalizable neural network with embeddings in $R^{2}$ for a mathematical reasoning task such as the sparse modular addition problem. We obtain similar conclusions than in Section A.2 with successful models having more dense activations.